# A Non-Parametric Sensitivity Analysis for Bounding Bias in Hybrid Control Trials

### v1

@@ 5 Discussion
The findings in the Section 4 imply that there needs to be careful deliberation in selecting the ideal ratio of RCT data against the external control data. From Figure 3, we see finite sample size issues with type I error inflation when data are heavily imbalanced, even when there is no bias present. This finding stems from the form of the second-order remainder of the trial-specific ATE as described in Appendix B.1 where it takes longer for the asymptotics to ‘kick in’ compared to the RCT ATE when $\frac{1}{\hat{q}}$ is small. We believe this to be an important finding, as the form of the remainder has not been explored in prior trial-specific ATE literature, despite its relevance to understanding the asymptotic behavior of estimators under certain conditions, such as varying values of $q$. However, regardless of the external sample sizes in the simulations, these issues seem to disappear once the RCT sample size reaches 100. As a result, we recommend being wary of using unbalanced ratios when RCT sample size is especially small. However, as seen in Figure 4, efficiency can be safely improved through the combined use of external data and sensitivity analysis even when using smaller external sample sizes.
@@
When RCT sample size is sufficiently large (100$\approx$ 100≈ 100), the possibility of using larger external sample sizes arises. The question of how large depends on several factors. First, we must consider the expected effect size of treatment, and then we must consider how different the external control arm may be from the RCT control arm. However, this second consideration should not be a major concern given proper selection of the external control arm. In fact, it would be improper to proceed with HCT analysis if the external control arm were expected to be significantly different than the RCT control arm. Despite this, it is worth considering the size of the inevitable bias from confounding since it can be seen in Figures 2 and 4 that any difference is amplified when external sample size is increased. The expected effect size is then important here; if the expected effect size is relatively small, amplified bias threatens correctly identifying a treatment effect. In this case, we recommend using a smaller external control arm. However, in the case that the expected effect size is relatively large (such as in the simulations), we see in Figure 4 that we can safely increase power by increasing the external sample size. The greatest power gains were achieved in the simulations when external sample size was 500 or 1000. While the gains in power are large from 100 to 500, the gains in power are relatively smaller from 500 to 1000. Consequently, it may be best to still limit the ratio of data away from extreme imbalances since the potential consequences outweigh the marginal benefits of relying heavily on external data.
@@
These considerations, however, all rely on the assumption made in the simulations that the true values of $C^{2}_{Y}$ and $C^{2}_{D}$ are known. In reality, the size of confounding due to omitted variable bias is never truly known. As mentioned above, omitted variable bias and sensitivity literature has explored different benchmarking algorithms to determine the relative plausible sizes of unobserved confounding [8, 14, 15, 21, 11]. In particular, benchmarking methods are popular where the observed covariates are used to approximate the influence of unobserved confounders. It has been argued that given strategic selection of measured covariate variables, the unobserved covariates likely are equally or less explanatory of outcome and treatment assignment [8]. With this assumption, one can use observed covariates to estimate bounds on the relationships of any unobserved variables with outcome and treatment. McClean et al. [32] summarizes these methods where notable strategies include the “leave-one-out” approach, addressed earlier in the methods section, as well as more complicated methods such as “leave-some-out” and “leave-every-out” to also rely on subsets of observed covariates. In all, these algorithmic approaches hold promise as benchmarking methods and can be used in combination with this sensitivity analysis to estimate plausible ranges of values for $C^{2}_{Y}$ and $C^{2}_{D}$ in order to judge robustness against omitted variable bias. While benchmarking itself provides measures of confounding due to unobserved confounders, critics have pointed out that relying solely on benchmarking may lead to erroneous conclusions [15, 32]. The benefit of using these benchmarks in conjunction with the sensitivity analysis outlined here is that the bias is properly identified, leading to safer conclusions on the causal effect of treatment [15].
@@
Finally, the scope of this paper focuses on binary outcomes for simplicity. However, we expect these findings to generalize easily to the continuous outcome scenario. The main motivation behind focusing on binary outcomes was the $r(X)$ term, the ratio of variance in the outcome between the RCT control data and external control data, in the Li estimator for the trial-specific average treatment effect. To reduce the number of values to be estimated, we binarized the outcome such that $r(X)$ would equal 1 under mean exchangeability. For continuous outcomes where $r(X)$ would need to be estimated, the original paper provides insight on empirical methods [28]. Further, we believe that the results would translate to continuous outcomes since the variance ratio is expected to only affect efficiency, not bias. Therefore, the idea of increasing trial efficiency while protecting against bias with the sensitivity analysis should be able to be extended to continuous outcomes after some verification with simulations.
@@
For estimation, we relied on an efficient estimating equations approach with debiased machine learning. However, so long as the framework in the paper is followed, other estimation techniques can also be used. For example, methods such as targeted maximum likelihood estimation (TMLE) may be applied after solving for the canonical gradient of the bias bound [14, 25]. We however relied on an EEE approach for an intuitively understandable approach by prospective researchers who may implement the sensitivity analysis. Further, we relied on the Li estimator for the trial-specific ATE due to its efficiency, but the sensitivity analysis can be applied for other estimators of the trial-specific ATE so long as the Riesz representer and outcome regression can be properly identified.
@@
For selection of the external control population, it is important that the external control subjects are as similar as possible to the RCT control subjects to maintain the integrity of results. Several guidelines exist for selection of external controls, such as Pocock’s selection criteria [34]. In particular, the external control subjects must have:
@@
These strict criteria attempt to ensure that the external controls are as similar as possible to the randomized control subjects. However, it is argued in some literature that these can be relaxed in extreme cases where data is limited or there is a pressing need to conduct research [30]. For example, there has been movement towards also including electronic health record (EHR) data as a source of external control data [12]. The implication of adding EHR data as an external data source is large, vastly enriching the availability of recent, relevant health data. Regardless of a relaxation of these criteria, data selection should still mirror these criteria as closely as practically possible to balance minimizing biased decision making and minimizing the patient burden.
@@
Other attempts to minimize bias from external control selection include a priori approaches such as a test-then-pool approach [42]. In this approach, a simple hypothesis test is conducted on the outcome prior to the decision to pool the external and randomized controls. Given an insignificant finding, researchers pool the data together and analyze the data as an HCT. Given a significant finding, the hypothesis test raises a red flag, resulting in researchers analyzing the RCT separately from the external data. Li et al. [28] also suggest a test for the mean exchangeability assumption in which a parametric outcome model is fit with interactions between $D$ and $X$. A variety of other tests exist in which this mean exchangeability assumption can be tested in advance of pooling such as the non-parametric omnibus test of equality [31] However, these approaches are all limited in that they are only asymptotically guaranteed to detect discrepancies in the data. Since there will always be a level of bias between the control groups, these tests will always choose to not pool the data asymptotically. However, these are applied to small samples in the HCT setting where the availability of data (from the RCT and/or external data source) is constrained. As a result, these will only identify any significantly large discrepancies in the data. The sensitivity analysis is thus a logical complement to such tests (and pooling) to detect the smaller inevitable differences in the data that lead to bias in naive HCT analysis.
@@
The ability to safely use external control data in hybrid control trials has multiple implications. HCTs have been proposed as early as 1976, but this trial design has typically been discouraged by major regulatory bodies [1]. Despite openness to the use of external data, the reliance on the mean exchangeability assumption, which practically is always violated to some degree, has hindered the wide acceptance of HCTs. HCT literature has predominantly focused on either creating new, more efficient estimators or relaxing the mean exchangeability assumption. Bayesian methodologies have been proposed, in which one can dynamically adjust how much information to borrow from the external data based on compatibility of the controls, but gains in efficiency are minimal when enforcing strict type I error control [24]. However, HCT literature has not explored what one can do when the mean exchangeability assumption is not assumed to be met. This paper fills in this gap in the research, quantifying this inevitable bias through a sensitivity analysis. We hope that filling in this gap in the research and making findings more credible will lead to the wider acceptance of HCTs.
@@
The benefits of the HCT trial design are numerous. Of course, there is reduced recruitment, leading to shorter timelines and budgets, but the benefits go beyond these factors. Ethically, the HCT trial design is attractive as it minimizes patient burden [30]. A study relying on solely an RCT requires more patients to be randomized, resulting in a larger number of patients missing out on receiving a treatment. In addition, with increased control information, it is possible to change the randomization ratio such that more RCT participants receive the treatment than the control. This not only is more ethical but can also be more attractive to prospective participants, encouraging trial participation. Another benefit of adding credibility to HCTs is that it can decrease reliance on single-arm study designs in which all control data come from external sources. Single-arm trials are also used in small sample settings, but are even more dependent on the mean exchangeability assumption, resulting in an even higher risk of bias [38]. While there are cases in which even an HCT may be infeasible, the cases in which an HCT is feasible should provide a safer alternative trial design than the single-arm trial [30].
@@
Considering its extreme nature, HCTs are only recommended in situations where there is a pressing need to borrow information from outside of an RCT. In particular, this includes rare disease research, oncology trials, and traditionally excluded populations from research. Due to the nature of these research areas, conducting a full-scale RCT is usually impractical, hindering the development of medicines and treatments for these populations. The findings of this paper align with the aims of the 21st Century Cures Act “to accelerate the discovery, development, and delivery of 21st century cures” [2]. With added credibility to the findings from HCTs, regulatory bodies will hopefully allow for more HCT trials, greatly increasing the volume of research that can be dedicated to these traditionally difficult areas in medical research.
@@ A.3 Riesz Representer of Trial-Specific ATE
$\quad\left.+\pi(X)E[Y|A=0,X,D=0]\frac{(1-\pi(X))r(X)}{\pi(X)(1-p(X))+(1-\pi(X))r(X)}\right]$ (2, 3)
@@
$=\frac{1}{q}E\left[\underbrace{\left(\frac{DA}{p(X)}-W(A,X,D)\right)}_{\text{Riesz Representer}}E[Y|A,X,D]\right]$
Assumptions: (1) Positivity, (2) Total Expectation, (3) Mean Exchangeability of Controls
Note that since we only use mean exchangeability for the control units, we also use our knowledge that $P(A=1|X,D=0)=0$ and $P(A=0|X,D=0)=1$.
@@ B.1 Second-Order Remainder of Trial-Specific ATE
Given the consistency of $\hat{\Psi}$ and the central limit theorem applied to $\hat{q}$, the first term is $o_{P}(n^{-1/2})$ as long as $q\neq 0$. The second term in the remainder is a typical “doubly robust” term: $\frac{1}{\hat{q}}E(\hat{\beta}(A,X,D)-\beta(A,X,D))(\hat{\mu}(A,X)-\mu(A,X))$. Typically for second-order convergence of the double robust term, we require that $\|\hat{\mu}-\mu\|\|\hat{\beta}-\beta\|$ is $o_{P}(n^{-\frac{1}{2}})$ [35]. In our setting we must also require $q$ bounded away from 0. This is not an issue.
@@
Appendix C Type I Error Rate under Two Repetition Cross Fitting

### v2

@@ 6 Discussion
The findings in Section 4 imply that there needs to be careful deliberation in selecting the ideal ratio of RCT data against the external control data. From Figure 3, we see finite sample size issues with type I error inflation when data are heavily imbalanced, even when there is no bias present. This finding stems from the form of the second-order remainder of the trial-specific ATE as described in Appendix B.1 where it takes longer for the asymptotics to ‘kick in’ compared to the RCT ATE when $\hat{q}$ is small. We believe this to be an important finding, as the form of the remainder has not been explored in prior trial-specific ATE literature, despite its relevance to understanding the asymptotic behavior of estimators under certain conditions, such as varying values of $q$. However, regardless of the external sample sizes in the simulations, these issues seem to disappear once the RCT sample size reaches 100. As a result, we recommend being wary of using unbalanced ratios when RCT sample size is especially small. However, as seen in Figure 4, efficiency can be safely improved through the combined use of external data and sensitivity analysis even when using smaller external sample sizes.
@@
When RCT sample size is sufficiently large, the possibility of using larger external sample sizes arises. The question of how large depends on several factors. First, we must consider the expected effect size of treatment, and then we must consider how different the external control arm may be from the RCT control arm. However, this second consideration should not be a major concern given proper selection of the external control arm. In fact, it would be improper to proceed with HCT analysis if the external control arm were expected to be significantly different than the RCT control arm. Despite this, it is worth considering the size of the inevitable bias from confounding since it can be seen in Figures 2 and 4 that any difference is amplified when external sample size is increased. The expected effect size is then important here; if the expected effect size is relatively small, amplified bias threatens correctly identifying a treatment effect. In this case, we recommend using a smaller external control arm. However, in the case that the expected effect size is relatively large (such as in the simulations), we see in Figure 4 that we can safely increase power by increasing the external sample size. The greatest power gains were achieved in the simulations when external sample size was 500 or 1000. While the gains in power are large from 100 to 500, the gains in power are relatively smaller from 500 to 1000. Consequently, it may be best to still limit the ratio of data away from extreme imbalances since the potential consequences outweigh the marginal benefits of relying heavily on external data.
@@
These considerations, however, all rely on the assumption made in the simulations that the true values of $C^{2}_{Y}$ and $C^{2}_{D}$ are known. In reality, the size of confounding due to omitted variable bias is never truly known. As mentioned above, omitted variable bias and sensitivity literature has explored different benchmarking algorithms to determine the relative plausible sizes of unobserved confounding [altonji_selection_2005, chernozhukov_long_2024, cinelli_making_2020, imbens_sensitivity_2003, bonvini_sensitivity_2022]. In particular, benchmarking methods are popular where the observed covariates are used to approximate the influence of unobserved confounders. It has been argued that given strategic selection of measured covariate variables, the unobserved covariates likely are equally or less explanatory of outcome and treatment assignment [altonji_selection_2005]. With this assumption, one can use observed covariates to estimate bounds on the relationships of any unobserved variables with outcome and treatment. mcclean_calibrated_2024 summarizes these methods where notable strategies include the “leave-one-out” approach, addressed earlier in the methods section, as well as more complicated methods such as “leave-some-out” and “leave-every-out” to also rely on subsets of observed covariates. In all, these algorithmic approaches hold promise as benchmarking methods and can be used in combination with this sensitivity analysis to estimate plausible ranges of values for $C^{2}_{Y}$ and $C^{2}_{D}$ in order to judge robustness against omitted variable bias. While benchmarking itself provides measures of confounding due to unobserved confounders, critics have pointed out that relying solely on benchmarking may lead to erroneous conclusions [cinelli_making_2020, mcclean_calibrated_2024]. The benefit of using these benchmarks in conjunction with the sensitivity analysis outlined here is that the bias bound is properly identified, leading to safer conclusions on the causal effect of treatment [cinelli_making_2020].
@@
Finally, the scope of this paper focuses on binary outcomes for simplicity. However, we expect these findings to generalize easily to the continuous outcome scenario. The main motivation behind focusing on binary outcomes was the $r(X)$ term, the ratio of variance in the outcome between the RCT control data and external control data, in the Li estimator for the trial-specific average treatment effect. To reduce the number of values to be estimated, we binarized the outcome such that $r(X)$ would equal 1 under mean exchangeability. For continuous outcomes where $r(X)$ would need to be estimated, the original paper provides insight on empirical methods [li_improving_2023]. Further, we believe that the results would translate to continuous outcomes since the variance ratio is expected to only affect efficiency, not bias. Therefore, the idea of increasing trial efficiency while protecting against bias with the sensitivity analysis should be able to be extended to continuous outcomes after some verification with simulations.
@@
For estimation, we relied on an efficient estimating equations approach with debiased machine learning. However, so long as the framework in the paper is followed, other estimation techniques can also be used. For example, methods such as targeted maximum likelihood estimation (TMLE), which also uses the efficient influence function, may be implemented [chernozhukov_long_2024, laan_targeted_2011]. However, we relied on an EEE approach for an intuitively understandable approach by prospective researchers who may implement the sensitivity analysis. Further, we relied on the Li estimator for the trial-specific ATE due to its efficiency, but the sensitivity analysis can be applied for other estimators of the trial-specific ATE so long as the Riesz representer and outcome regression can be properly identified.
@@
For selection of the external control population, it is important that the external control subjects are as similar as possible to the RCT control subjects to maintain the integrity of results. Several guidelines exist for selection of external controls, such as Pocock’s selection criteria [pocock_combination_1976]. In particular, the external control subjects must have:
@@
These strict criteria attempt to ensure that the external controls are as similar as possible to the randomized control subjects. However, it is argued in some literature that these can be relaxed in extreme cases where data is limited or there is a pressing need to conduct research [lim_minimizing_2018]. For example, there has been movement towards also including electronic health record (EHR) data as a source of external control data [carrigan]. The implication of adding EHR data as an external data source is large, vastly enriching the availability of recent, relevant health data. Regardless of a relaxation of these criteria, data selection should still mirror these criteria as closely as practically possible to balance minimizing biased decision making and minimizing the patient burden. Further, as discussed in Section 5, careful selection of the external control sample can help loosely control the magnitude of $C^{2}_{D}$. When population differences are smaller, the potential impact a confounder has on trial participation is likewise smaller. Although this method works even if the populations are dissimilar, it is important that researchers carefully select external control samples to avoid inducing unnecessary bias in estimation.
@@
Other attempts to minimize bias from external control selection include a priori approaches such as a test-then-pool approach [viele_use_2014]. In this approach, a simple hypothesis test is conducted on the outcome prior to the decision to pool the external and randomized controls. Given an insignificant finding, researchers pool the data together and analyze the data as an HCT. Given a significant finding, the hypothesis test raises a red flag, resulting in researchers analyzing the RCT separately from the external data. li_improving_2023 also suggest a test for the mean exchangeability assumption in which a parametric outcome model is fit with interactions between $D$ and $X$. A variety of other tests exist in which this mean exchangeability assumption can be tested in advance of pooling such as the non-parametric omnibus test of equality [luedtke_omnibus_2019] However, these approaches are all limited in that they are only asymptotically guaranteed to detect discrepancies in the data. Since there will always be a level of bias between the control groups, these tests will always choose to not pool the data asymptotically. However, these are applied to small samples in the HCT setting where the availability of data (from the RCT and/or external data source) is constrained. As a result, these will only identify any significantly large discrepancies in the data. The sensitivity analysis is thus a logical complement to such tests (and pooling) to detect the smaller inevitable differences in the data that lead to bias in naive HCT analysis.
@@
The ability to safely use external control data in hybrid control trials has multiple implications. HCTs have been proposed as early as 1976, but this trial design has typically been discouraged by major regulatory bodies [noauthor_international_2000]. Despite openness to the use of external data, the reliance on the mean exchangeability assumption, which practically is always violated to some degree, has hindered the wide acceptance of HCTs. HCT literature has predominantly focused on either creating new, more efficient estimators or relaxing the mean exchangeability assumption. For example, Bayesian methodologies have been proposed, in which one can dynamically adjust how much information to borrow from the external data based on compatibility of the controls, but gains in efficiency are minimal when enforcing strict type I error control [koppschneider_power_2020]. Previous HCT literature focuses on mitigating biases through dynamic methods whereas we focus on quantifying bias under the worst case scenario. This paper fills in this gap in the research, quantifying this inevitable bias through a sensitivity analysis. We hope that filling in this gap in the research and making findings more credible will lead to the wider acceptance of HCTs.
@@
The benefits of the HCT trial design are numerous. Of course, there is reduced recruitment, leading to shorter timelines and budgets, but the benefits go beyond these factors. Ethically, the HCT trial design is attractive as it minimizes patient burden [lim_minimizing_2018]. A study relying on solely an RCT requires more patients to be randomized, resulting in a larger number of patients missing out on receiving a treatment. In addition, with increased control information, it is possible to change the randomization ratio such that more RCT participants receive the treatment than the control. This not only is more ethical but can also be more attractive to prospective participants, encouraging trial participation. Another benefit of adding credibility to HCTs is that it can decrease reliance on single-arm study designs in which all control data come from external sources, where possible. Single-arm trials are also used in small sample settings, but rely on a stronger version of the mean exchangeability assumption, resulting in an even higher risk of bias [tan_augmenting_2022]. While there are cases in which even an HCT may be infeasible, the cases in which an HCT is feasible should provide a safer alternative trial design than the single-arm trial [lim_minimizing_2018].
@@
Considering its extreme nature, HCTs are only recommended in situations where there is a pressing need to borrow information from outside of an RCT. In particular, this includes rare disease research, oncology trials, and traditionally excluded populations from research. Due to the nature of these research areas, conducting a full-scale RCT is usually impractical, hindering the development of medicines and treatments for these populations. The findings of this paper align with the aims of the 21st Century Cures Act “to accelerate the discovery, development, and delivery of 21st century cures” [noauthor_21st_2017]. With added credibility to the findings from HCTs, regulatory bodies will hopefully allow for more HCT trials, greatly increasing the volume of research that can be dedicated to these traditionally difficult areas in medical research.
[…]
@@ A.3 Riesz Representer of Trial-Specific ATE
$\quad\left.+\pi(X)E[Y|A=0,X,D=0]\frac{(1-\pi(X))r(X)}{\pi(X)(1-p(X))+(1-\pi(X))r(X)}\right]$ (2)
@@
$=E\left[\underbrace{\frac{1}{q}\left(\frac{DA}{p(X)}-W(A,X,D)\right)}_{\text{Riesz Representer}}E[Y|A,X,D]\right]$
Assumptions: (1) Positivity, (2) Mean Exchangeability of Controls
Note that since we only use external data for control data, we also use our knowledge that $P(A=1|X,D=0)=0$ and $P(A=0|X,D=0)=1$ to construct weights. Further, in the second expansion of $q[E[E[Y|A=0,X]]$ we use optimal efficiency weights derived from the influence function of the target parameter (which combine propensity score weights and the variance ratio $r(X)$).
@@ B.1 Second-Order Remainder of Trial-Specific ATE
Given the consistency of $\hat{\Psi}$ and the central limit theorem applied to $\hat{q}$, the first term is $o_{P}(n^{-1/2})$ as long as $q\neq 0$. The second term in the remainder is a typical “doubly robust” term: $\frac{1}{\hat{q}}E(\hat{\beta}(A,X,D)-\beta(A,X,D))(\hat{\mu}(A,X)-\mu(A,X))$. Typically for second-order convergence of the double robust term, we require that $\|\hat{\mu}-\mu\|\|\hat{\beta}-\beta\|$ is $o_{P}(n^{-\frac{1}{2}})$ [intro]. In our setting we must also require $q$ bounded away from 0. This is not an issue.
@@
Appendix C Additional Figures
C.1 Type I Error Rate under Two Repetition Cross Fitting
@@
C.2 Contour Plot for Case Study under $\rho=1$
Figure 7: Sensitivity contour plot for the HCT under $\rho=1$ for the most conservative estimates. The unadjusted estimate of the lower bound of the trial-specific average treatment effect is reported. The red dotted line represents the point at which results are no longer statistically significant under combinations of $C^{2}_{Y}$ and $C^{2}_{D}$.